OpenStudy (anonymous):

solve the linear system: 8x-3y=3 3x-2y+5=0

OpenStudy (anonymous):

how did you get that?

jhonyy9 (jhonyy9):

8x-3y=3 / multiply by 2 3x-2y=-5 / multiply by -3 16x-6y=6 -9x+6y=15 -------------- assum 7x 0 = 21 7x=21 x=21/7 x=3 3x-2y=-5 3*3-2y=-5 9+5=2y 2y=14 y=14/2 y=7 x=3 y=7

jhonyy9 (jhonyy9):

hope understandably sure

jhonyy9 (jhonyy9):

good luck bye